# Survey on Publicly Available Sinhala Natural Language Processing Tools and Research

### v21

@@ Abstract
Sinhala is the native language of the Sinhalese people who make up the largest ethnic group of Sri Lanka. The language belongs to the globe-spanning language tree, Indo-European. However, due to poverty in both linguistic and economic capital, Sinhala, in the perspective of Natural Language Processing tools and research, remains a resource-poor language which has neither the economic drive its cousin English has nor the sheer push of the law of numbers a language such as Chinese has. A number of research groups from Sri Lanka have noticed this dearth and the resultant dire need for proper tools and research for Sinhala natural language processing. However, due to various reasons, these attempts seem to lack coordination and awareness of each other. The objective of this paper is to fill that gap of a comprehensive literature survey of the publicly available Sinhala natural language tools and research so that the researchers working in this field can better utilize contributions of their peers. As such, we shall be uploading this paper to arXiv and perpetually update it periodically to reflect the advances made in the field.
@@ 1 Introduction
Sinhala [^1] Englebretson and Genetti [1] observe that in some contexts the Sinhala language is also referred as Sinhalese, Singhala, and Singhalese language, being the native language of the Sinhalese people [2, 3, 4], who make up the largest ethnic group of the island country of Sri Lanka, enjoys being reported as the mother tongue (L1) of approximately 16 million people [5, 6]. When both L1 and L2 speakers are counted, $79.7\%$ of the total Sri Lankan population are literate in Sinhala [7]. To give a brief linguistic background for the purpose of aligning the Sinhala language with the baseline of English, primarily it should be noted that Sinhala language belongs to the same the Indo-European language tree [8, 9]. However, unlike English, which is part of the Germanic branch, Sinhala belongs to the Indo-Aryan branch. Further, Sinhala, unlike English, which borrowed the Latin alphabet, has its own writing system, which is a descendant of the Indian Brahmi script [10, 11, 12, 13, 14, 15, 16]. By extension, this makes Sinhala Script a member of the Aramaic family of scripts [17, 18]. Thus by inheritance, Sinhala writing system is abugida (alphasyllabary), which to say, that consonant-vowel sequences are written as single units [19]. We show the evolutionary eras of the Sinhala script in Appendix A. It should be noted that the modern Sinhala language have loanwords from languages such as Tamil, English, Portuguese, and Dutch due to various historical reasons [20]. Regardless of the rich historical array of literature spanning several millennia (starting between $3^{rd}$ to $2^{nd}$ century BCE [21, 22]), modern natural language processing tools for the Sinhala language are scarce [23].
@@
Figure 1: NLP layers and tasks [24]
@@
The objective of this paper is to serve as a comprehensive survey on the state of natural language processing resources for the Sinhala language. The initial structure and content of this survey are heavily influenced by the preliminary surveys carried out by de Silva [23] and Wijeratne et al. [24]. However, our hope is to host this survey at arXiv as a perpetually evolving, living research article [34, 35] which continuously gets updated as new research and tools for Sinhala language are created and made publicly available. We also discuss how the non-compliance of policies to put data and code online [36], after the research is concluded and the paper is published, have negatively impacted the growth and sustainability of Sinhala NLP. Hence, it is our hope that this work will help future researchers who are engaged in Sinhala NLP research to conduct their literature surveys efficiently and comprehensively. For the success of this survey, we shall also consider the Sri Lankan NLP tools repository, lknlp [^3] https://github.com/lknlp/lknlp.github.io. This manuscript is at version $5.3.0$. The latest version of the manuscript can be obtained from arXiv [^4] https://arxiv.org/abs/1906.02358 or ResearchGate [^5] http://bit.ly/31AhvvR.
@@
Figure 6 in Appendix B shows the most prolific researchers in the domain of Sinhala NLP. The nodes contain the name of the researcher along with the total number of Sinhala NLP papers that researcher has authored. The edges between two researchers are labeled with the number of Sinhala NLP papers the relevant pair of researchers have co-authored. When selecting authors, we have applied a threshold of 3 Sinhala NLP publications. Given that the objective of the visualization is to portray the co-operation between researchers, we have also added the strongest edge that connects each researcher to the rest of the researchers in the graph. The few isolated nodes are researchers that have authored at least 3 Sinhala NLP publications but do not have any coauthored papers with anyone else on the graph. We have also added labels to clusters in cases where all or the majority of researchers in those clusters have the same affiliation. It is observable that the cluster from the Department of Computer Science & Engineering, University of Moratuwa is the most prolific in Sinhala NLP research.
Figure 7 in Appendix B shows the probability of studies from the institutions to which the most prolific authors from Figure 6 are affiliated citing institutions of the same set. We observe both interesting and disturbing trends which we discuss in the figure caption
@@
Figure 8 in Appendix B shows the mapping between authors with at least 10 papers in the Sinhala NLP domain and their research interests denoted by the subsection titles of the Section 3 of this paper. If two or more authors on the diagram have co-authored a paper, it gets counted for each of the authors separately without any bias on the author order listed on the publication. For example, the paper Building a wordnet for Sinhala [37] is counted for both Nisansa de Silva and Gihan Dias. If a single paper contributes to more than one research area, that paper is counted for all of the research areas to which it contributes. For example, the paper Sinhala Text Classification: Observations from the Perspective of a Resource Poor Language [23] which introduces a new Sinhala text classification data set is counted for both Data Sets (Section 3.2) and Text Classification (Section 3.10).
@@ 3.1 Corpora
For any language, the key for NLP applications and implementations is the existence of adequate corpora. On this matter, a relatively substantial Sinhala text corpus [^9] https://osf.io/a5quv/ was created by Upeksha et al. [86, 87] by web crawling. It was later extended by adding Jathaka Stories [^10] https://bit.ly/JathakaTxt and more web-crawled news articles [^11] https://bit.ly/3osodBj. Later a smaller Sinhala news corpus [^12] https://osf.io/tdb84/ was created by de Silva [23]. Both of the above corpora are publicly available. However, none of these come close to the massive capacity and range of the existing English corpora. A word corpus of approximately 35,000 entries was developed by Weerasinghe et al. [88]. But it does not seem to be online anymore. Guzmán et al. [89] provided two monolingual corpora for Sinhala. Those were a 155k+ sentences of filtered Sinhala Wikipedia [^13] http://bit.ly/2EQZ7oM and 5178k+ sentences of Sinhala common crawl [^14] http://bit.ly/2ZaQFZo.  Wijeratne and de Silva [90] have publicly released [^15] https://bit.ly/2GEI4d6 a massive corpus of text and stop words taken from a decade of Sinhala Facebook posts. While the stop word extraction of Wijeratne and de Silva [90] algorithmic, Lakmal et al. [91] has introduced a manually curated stop word list [^16] http://bit.ly/3wFi0Wf. A parallel corpus of Sinhala and English was collected by Bañón et al. [92] containing 217,407 sentences and available to download from their website [^17] https://www.paracrawl.eu/. However, the later audit by Caswell et al. [93] raised issues on the quality of that data set. A parallel corpus [^18] https://github.com/kdissa/comparable-corpus of aligned Sinhala-English documents and sentences obtained from crawling the web was released by Sachintha et al. [94]. The study by Warusawithana et al. [95] created a refined version [^19] https://bit.ly/3JF77t9 of the OpenSLR-52 speech corpus for Sinhala [^20] https://openslr.org/52 by Kjartansson et al. [96] [^21] A later further overview on the corpus is also available [97].Dhananjaya et al. [98] created sin-cc-15M corpus2222223https://tinyurl.com/42un7a9y, which they claim, at the time of their publication to be the largest monolingual Sinhala corpus. A text corpus collected from the Sihala blog Kalaya [^23] http://www1.kalaya.org/ along with the relevant code is available on Github [^24] https://bit.ly/KalayaCorpus.
@@ 3.2 Data Sets
Facebook has released FastText [109, 110, 111] models for the Sinhala language trained using the Wikipedia corpus. They are available as both text models [^30] http://bit.ly/2JXAyL8 and binary files [^31] http://bit.ly/2JY5J9c. Using the above models by Facebook, Lakmal et al. [91] have created an extended FastText model trained on Wikipedia, News, and official government documents. The binary file [^32] http://bit.ly/2WowH0h of the trained model is available to be downloaded. Herath et al. [112], Herath and Medagoda [113] have compiled a report on the Sinhala lexicon for the purpose of establishing a basis for NLP applications. A comparative analysis of Sinhala word embedding has been conducted by Lakmal et al. [91]. A similar study was then conducted by Silva [114] analyising the progress that had been made in NLP in the intervening years. A dataset [^33] https://github.com/LIRNEasia/MisinformationCorpusSinhala consisting of 3576 Sinhala documents drawn from Sri Lankan news websites and tagged (CREDIBLE, FALSE, PARTIAL or UNCERTAIN) was published by Jayawickrama et al. [115] A benchmark data set for Sinhala spell correction was created by  Sonnadara et al. [116] which they put online on Github [^34] https://github.com/chason94/SinNeuSpellCorrector. De Saa and Ranathunga [117] has released a data set [^35] https://bit.ly/3FTNaMZ for Sinhala hate speech detection which consists of comments from youtube. Another data set consisting of Sinhala hate speech comments pulled from Facebook is available [^36] https://bit.ly/3j9r3Za on Kaggle but it does not have an accompanying paper. Similarly, Perera et al. [118] have released a hate annotated dataset [^37] https://github.com/Isurie/Text-Classification-Module/tree/master/Dataset of 1600+ annotated Sinhala tweets. A text-to-speech data set [^38] https://github.com/pnfo/sinhala-tts-dataset with 3300 Sinhala sentences with 7.5 hours of recordings was released by the Path Nirvana Foundation. The open parallel corpus (OPUS) [^39] https://opus.nlpl.eu/ by Tiedemann [119] curates the largest collection of parallel data between Sinhala and other languages. Zhang et al. [120] sampled data from OPUS for 100 languages (including Sinhala) and created the OPUS100404040https://github.com/EdinburghNLP/opus-100-corpus data set as a benchmark for neural machine translation (NMT). Thier code is also available on GitHub [^41] https://github.com/bzhangGo/zero. Further, Tiedemann [121] created an NMT benchmark named Tatoeba MT Challenge [^42] https://github.com/Helsinki-NLP/Tatoeba-Challenge using OPUS data based on Tatoeba [^43] https://tatoeba.org/. The English-Sinhala data set can also be directly downloaded from a link on Github [^44] https://github.com/Helsinki-NLP/Tatoeba-Challenge/tree/master/data.
@@
Ranasinghe et al. [128] created a benchmark dataset they named SOLD [^63] https://huggingface.co/datasets/sinhala-nlp/SOLD: Sinhala Offensive Language Dataset which contains 10,000 posts from Twitter annotated both at sentence-level and token-level with the two classes offensive and not offensive. In the same paper, they also introduce the dataset SemiSOLD [^64] https://huggingface.co/datasets/sinhala-nlp/SemiSOLD which contains 145,000 Sinhala tweets annotated with the same classes but with a semi-supervised approach. The relevant code is also available on Github  [^65] https://github.com/Sinhala-NLP/SOLD. Sinhala is included in the 2800+ language metadata set composed by van Esch et al. [129]. This can be used for comparative analysis of Sinhala against other languages in the data set. Ruder et al. [130] created the multi-language dataset named XTREME-UP [^66] https://github.com/google-research/xtreme-up which contains Sinhala data sets for OCR and Transliteration tasks. Pratap et al. [131] created the The Massively Multilingual Speech (MMS) [^67] https://github.com/facebookresearch/fairseq/tree/main/examples/mms data set which contains Sinhala for the Spoken Language Identification (LID) task. A data set [^68] https://github.com/laurieburchell/open-lid-dataset and a model for the Language Identification (LID) task including Sinhala were created by Burchell et al. [132]. A large Sinhala-English dictionary with $1,368,416$ unfiltered and $195,255$ filtered En-Sipairs has been made publicly available on github [^69] https://github.com/kasunw22/sinhala-para-dict/tree/main by Wickramasinghe and De Silva [133]. In their follow-up work [134] they used that dictionary data to obtain Sinhala-English embedding alignment. The relevant code and data can be accessed on GitHub [^70] https://bit.ly/3t3SKu7. Nguyen et al. [135] released CulturaX [^71] https://huggingface.co/datasets/uonlp/CulturaX a cleaned dataset for training Large Language Models. It contains 753,655 Sinhala documents with 880,289,097 tokens. However, it should be noted that this only amounts to 0.01% of their multilingual data set. In comparison, Tamil has a 0.07% share, Hindi has a 0.27% share, and English has a 45.13% share. Kudugunta et al. [136] released MADLAD-400727272https://github.com/google-research/google-research/tree/master/madlad_400 a manually audited, general domain, large document level dataset spanning 419 languages. For Sinhala, they report that they have 788K noisy docs, 349.2K clean docs, 22.1M noisy sentences, 16M clean sentences, 3.4B noisy tokens, and 1.9B clean tokens. Further, for the translation tasks they report that they have 7363378 sentence pairs where one of the languages in each pair is Sinhala.

### v22

@@ Abstract
Sinhala is the native language of the Sinhalese people who make up the largest ethnic group of Sri Lanka. The language belongs to the globe-spanning language tree, Indo-European. However, due to poverty in both linguistic and economic capital, Sinhala, in the perspective of Natural Language Processing tools and research, remains a resource-poor language which has neither the economic drive its cousin English has nor the sheer push of the law of numbers a language such as Chinese has. A number of research groups from Sri Lanka have noticed this dearth and the resultant dire need for proper tools and research for Sinhala natural language processing. However, due to various reasons, these attempts seem to lack coordination and awareness of each other. The objective of this paper is to fill that gap of a comprehensive literature survey of the publicly available Sinhala natural language tools and research so that the researchers working in this field can better utilize the contributions of their peers. As such, we shall be uploading this paper to arXiv and perpetually update it periodically as a living research article to reflect the advances made in the field. This manuscript is at version $5.4.0$.
@@ 1 Introduction
Sinhala [^1] Englebretson and Genetti [1] observe that in some contexts the Sinhala language is also referred as Sinhalese, Singhala, and Singhalese language, being the native language of the Sinhalese people [2, 3, 4], who make up the largest ethnic group of the island country of Sri Lanka, enjoys being reported as the mother tongue (L1) of approximately 16 million people [5, 6]. When both L1 and L2 speakers are counted, $79.7\%$ of the total Sri Lankan population are literate in Sinhala [7]. A strong correlation between the island of Sri Lanka and the usage of Sinhala has been observed [8] [^2] Dunn and Edwards-Brown [8] in their work to augment Language Identification (LID) task with geography observe that for Sinhala LID does not benefit from the geography model given that the to models already has an alignment of 99%.. This implies two things: 1) the majority of the Sinhala linguistic sources are located in the geographical area of Sri Lanka, and 2) the majority of the available linguistic sources in Sri Lanka are in Sinhala.
To give a brief linguistic background for the purpose of aligning the Sinhala language with the baseline of English, primarily it should be noted that the Sinhala language belongs to the same Indo-European language tree [9, 10]. However, unlike English, which is part of the Germanic branch, Sinhala belongs to the Indo-Aryan branch. Further, Sinhala, unlike English, which borrowed the Latin alphabet, has its own writing system, which is a descendant of the Indian Brahmi script [11, 12, 13, 14, 15, 16, 17]. By extension, this makes Sinhala Script a member of the Aramaic family of scripts [18, 19]. Thus by inheritance, the Sinhala writing system is abugida (alphasyllabary), which to say, that consonant-vowel sequences are written as single units [20]. We show the evolutionary eras of the Sinhala script in Appendix A. It should be noted that the modern Sinhala language has loanwords from languages such as Tamil, English, Portuguese, and Dutch due to various historical reasons [21]. Regardless of the rich historical array of literature spanning several millennia (starting between $3^{rd}$ to $2^{nd}$ century BCE [22, 23]), modern natural language processing tools for the Sinhala language are scarce [24, 25]. Further, Sinhala is a diglossic language [26] which has two variations: the literary form and the spoken form.
@@
Figure 1: Traditional NLP layers and tasks (Derived from the work of Wijeratne et al. [27] and extended to encapsulate a wider range by this work.)
@@
The objective of this paper is to serve as a comprehensive survey on the state of natural language processing resources for the Sinhala language. The initial structure and content of this survey are heavily influenced by the preliminary surveys carried out by de Silva [24] and Wijeratne et al. [27]. However, our hope is to host this survey at arXiv as a perpetually evolving, living research article [37, 38] which continuously gets updated as new research and tools for Sinhala language are created and made publicly available. We also discuss how the non-compliance of policies to put data and code online [39], after the research is concluded and the paper is published, has negatively impacted the growth and sustainability of Sinhala NLP. Hence, it is our hope that this work will help future researchers who are engaged in Sinhala NLP research to conduct their literature surveys efficiently and comprehensively. For the success of this survey, we shall also consider the Sri Lankan NLP tools repository, lknlp [^4] https://github.com/lknlp/lknlp.github.io. This manuscript is at version $5.4.0$. The latest version of the manuscript can be obtained from arXiv [^5] https://arxiv.org/abs/1906.02358 or ResearchGate [^6] http://bit.ly/31AhvvR.
@@
Figure 6 in Appendix B shows the most prolific researchers in the domain of Sinhala NLP. The nodes contain the name of the researcher along with the total number of Sinhala NLP papers that the researcher has authored. The edges between the two researchers are labelled with the number of Sinhala NLP papers the relevant pair of researchers have co-authored. When selecting authors, we have applied a threshold of 3 Sinhala NLP publications. Given that the objective of the visualization is to portray the cooperation between researchers, we have also added the strongest edge that connects each researcher to the rest of the researchers in the graph. The few isolated nodes are researchers who have authored at least 3 Sinhala NLP publications but do not have any coauthored papers with anyone else on the graph. We have also added labels to clusters in cases where all or the majority of researchers in those clusters have the same affiliation. It is observable that the cluster from the Department of Computer Science & Engineering, University of Moratuwa is the most prolific in Sinhala NLP research.
@@
Figure 7 in Appendix B shows the probability of studies from the institutions to which the most prolific authors from Figure 6 are affiliated citing institutions of the same set. We observe both interesting and disturbing trends which we discuss in the figure caption. Figure 8 in Appendix B shows the mapping between authors with at least 10 papers in the Sinhala NLP domain and their research interests denoted by the subsection titles of the Section 3 of this paper. If two or more authors on the diagram have co-authored a paper, it gets counted for each of the authors separately without any bias on the author order listed on the publication. For example, the paper Building a wordnet for Sinhala [40] is counted for both Nisansa de Silva and Gihan Dias. If a single paper contributes to more than one research area, that paper is counted for all of the research areas to which it contributes. For example, the paper Sinhala Text Classification: Observations from the Perspective of a Resource Poor Language [24] which introduces a new Sinhala text classification data set is counted for both Data Sets (Section 3.2) and Text Classification (Section 3.10).
@@ 3.1 Corpora
For any language, the key for NLP applications and implementations is the existence of adequate corpora. On this matter, a relatively substantial Sinhala text corpus [^10] https://osf.io/a5quv/ was created by Upeksha et al. [89, 90] by web crawling. It was later extended by adding Jathaka Stories [^11] https://bit.ly/JathakaTxt and more web-crawled news articles [^12] https://bit.ly/3osodBj. Later a smaller Sinhala news corpus [^13] https://osf.io/tdb84/ was created by de Silva [24]. Both of the above corpora are publicly available. However, none of these come close to the massive capacity and range of the existing English corpora.
A word corpus of approximately 35,000 entries was developed by Weerasinghe et al. [91]. But it does not seem to be online anymore. Guzmán et al. [92] provided two monolingual corpora for Sinhala. Those were a 155k+ sentences of filtered Sinhala Wikipedia [^14] http://bit.ly/2EQZ7oM and 5178k+ sentences of Sinhala common crawl [^15] http://bit.ly/2ZaQFZo.  Wijeratne and de Silva [93] have publicly released [^16] https://bit.ly/2GEI4d6 a massive corpus of text and stop words taken from a decade of Sinhala Facebook posts. While the stop word extraction of Wijeratne and de Silva [93] algorithmic, Lakmal et al. [94] has introduced a manually curated stop word list [^17] http://bit.ly/3wFi0Wf. A parallel corpus of Sinhala and English was collected by Bañón et al. [95] containing 217,407 sentences and available to download from their website [^18] https://www.paracrawl.eu/. However, the later audit by Caswell et al. [96] raised issues on the quality of that data set. A parallel corpus [^19] https://github.com/kdissa/comparable-corpus of aligned Sinhala-English documents and sentences obtained from crawling the web was released by Sachintha et al. [97]. The study by Warusawithana et al. [98] created a refined version [^20] https://bit.ly/3JF77t9 of the OpenSLR-52 speech corpus for Sinhala [^21] https://openslr.org/52 by Kjartansson et al. [99] [^22] A later further overview on the corpus is also available [100].Dhananjaya et al. [101] created sin-cc-15M corpus2323233https://tinyurl.com/42un7a9y, which they claim, at the time of their publication to be the largest monolingual Sinhala corpus. A text corpus collected from the Sihala blog Kalaya [^24] http://www1.kalaya.org/ along with the relevant code is available on Github [^25] https://bit.ly/KalayaCorpus. A Sinhala NEWS corpus of more than 500,000 articles collected from various online free news websites was released [^26] https://github.com/Sinhala-NLP/NSINA by Hettiarachchi et al. [102].
@@ 3.2 Data Sets
Facebook has released FastText [113, 114, 115] models for the Sinhala language trained using the Wikipedia corpus. They are available as both text models [^32] http://bit.ly/2JXAyL8 and binary files [^33] http://bit.ly/2JY5J9c. Using the above models by Facebook, Lakmal et al. [94] have created an extended FastText model trained on Wikipedia, News, and official government documents. The binary file [^34] http://bit.ly/2WowH0h of the trained model is available to be downloaded. Herath et al. [116], Herath and Medagoda [117] have compiled a report on the Sinhala lexicon for the purpose of establishing a basis for NLP applications. A comparative analysis of Sinhala word embedding has been conducted by Lakmal et al. [94]. A similar study was then conducted by Silva [118] analyising the progress that had been made in NLP in the intervening years. A dataset [^35] https://github.com/LIRNEasia/MisinformationCorpusSinhala consisting of 3576 Sinhala documents drawn from Sri Lankan news websites and tagged (CREDIBLE, FALSE, PARTIAL or UNCERTAIN) was published by Jayawickrama et al. [119] A benchmark data set for Sinhala spell correction was created by  Sonnadara et al. [120] which they put online on Github [^36] https://github.com/chason94/SinNeuSpellCorrector. De Saa and Ranathunga [121] has released a data set [^37] https://bit.ly/3FTNaMZ for Sinhala hate speech detection which consists of comments from youtube. Another data set consisting of Sinhala hate speech comments pulled from Facebook is available [^38] https://bit.ly/3j9r3Za on Kaggle but it does not have an accompanying paper. Similarly, Perera et al. [122] have released a hate annotated dataset [^39] https://github.com/Isurie/Text-Classification-Module/tree/master/Dataset of 1600+ annotated Sinhala tweets. A text-to-speech data set [^40] https://github.com/pnfo/sinhala-tts-dataset with 3300 Sinhala sentences with 7.5 hours of recordings was released by the Path Nirvana Foundation. The open parallel corpus (OPUS) [^41] https://opus.nlpl.eu/ by Tiedemann [123] curates the largest collection of parallel data between Sinhala and other languages. Zhang et al. [124] sampled data from OPUS for 100 languages (including Sinhala) and created the OPUS100424242https://github.com/EdinburghNLP/opus-100-corpus data set as a benchmark for neural machine translation (NMT). Thier code is also available on GitHub [^43] https://github.com/bzhangGo/zero. Further, Tiedemann [125] created an NMT benchmark named Tatoeba MT Challenge [^44] https://github.com/Helsinki-NLP/Tatoeba-Challenge using OPUS data based on Tatoeba [^45] https://tatoeba.org/. The English-Sinhala data set can also be directly downloaded from a link on Github [^46] https://github.com/Helsinki-NLP/Tatoeba-Challenge/tree/master/data.
@@
Ranasinghe et al. [132] created a benchmark dataset they named SOLD [^65] https://huggingface.co/datasets/sinhala-nlp/SOLD: Sinhala Offensive Language Dataset which contains 10,000 posts from Twitter annotated both at sentence-level and token-level with the two classes offensive and not offensive. In the same paper, they also introduce the dataset SemiSOLD [^66] https://huggingface.co/datasets/sinhala-nlp/SemiSOLD which contains 145,000 Sinhala tweets annotated with the same classes but with a semi-supervised approach. The relevant code is also available on Github  [^67] https://github.com/Sinhala-NLP/SOLD. Sinhala is included in the 2800+ language metadata set composed by van Esch et al. [133]. This can be used for comparative analysis of Sinhala against other languages in the data set. Ruder et al. [134] created the multi-language dataset named XTREME-UP [^68] https://github.com/google-research/xtreme-up which contains Sinhala data sets for OCR and Transliteration tasks. Pratap et al. [135] created the The Massively Multilingual Speech (MMS) [^69] https://github.com/facebookresearch/fairseq/tree/main/examples/mms data set which contains Sinhala for the Spoken Language Identification (LID) task. A data set [^70] https://github.com/laurieburchell/open-lid-dataset and a model for the Language Identification (LID) task including Sinhala were created by Burchell et al. [136]. A large Sinhala-English dictionary with $1,368,416$ unfiltered and $195,255$ filtered En-Sipairs has been made publicly available on github [^71] https://github.com/kasunw22/sinhala-para-dict/tree/main by Wickramasinghe and De Silva [137]. In their follow-up work [138] they used that dictionary data to obtain Sinhala-English embedding alignment. The relevant code and data can be accessed on GitHub [^72] https://bit.ly/3t3SKu7. Nguyen et al. [139] released CulturaX [^73] https://huggingface.co/datasets/uonlp/CulturaX a cleaned dataset for training Large Language Models. It contains 753,655 Sinhala documents with 880,289,097 tokens. However, it should be noted that this only amounts to 0.01% of their multilingual data set. In comparison, Tamil has a 0.07% share, Hindi has a 0.27% share, and English has a 45.13% share. Kudugunta et al. [140] released MADLAD-400747474https://github.com/google-research/google-research/tree/master/madlad_400 a manually audited, general domain, large document level dataset spanning 419 languages. For Sinhala, they report that they have 788K noisy docs, 349.2K clean docs, 22.1M noisy sentences, 16M clean sentences, 3.4B noisy tokens, and 1.9B clean tokens. Further, for the translation tasks they report that they have 7363378 sentence pairs where one of the languages in each pair is Sinhala.
@@
The Dakshina dataset [^79] https://github.com/google-research-datasets/dakshina by [146] contains Native Sinhala script data (200k train and 28k validation sentences) along with Romanized (Singlish) data (10k sentences) from Wikipedia. It also contains a Romanization Lexicon (25k train and 5k test). Ranathunga et al. [147] have publicly released [^80] https://github.com/nlpcuom/quality-matters their data, code, and models for Sinhala-English translation. Very importantly, this includes a human-cleaned portion of the NLLB dataset [127] in En-Si [^81] https://huggingface.co/datasets/NLPC-UOM/nllb-top25k-ensi-cleaned.
Singh et al. [148] introduced the Aya Data set [^82] https://huggingface.co/datasets/CohereForAI/aya_dataset, a human-curated multilingual instruction fine-tuning (IFT) data set for large language models (LLM). They also introduced Aya Collection [^83] https://huggingface.co/datasets/CohereForAI/aya_collection, 44 instruction-style datasets that were created by transforming existing NLP datasets into pairs of prompt and completion and then using NMT on them to obtain data in multiple languages. Both of these data sets include Sinhala. Further, they also released Aya Evaluation suite [^84] https://huggingface.co/datasets/CohereForAI/aya_evaluation_suite to measure multilingual open-ended generation quality.